Every time Casey is at bat he has a 0.2 probability of getting on base (assume each at bat is an independent event, and that this probability never changes). In one week of baseball, he has twelve times at bat. What is the standard deviation of the number of times he will get on base in a week? [again, the On-Base Probability here might be different from the value given in the previous question] [must be accurate to 4 decimal places]
Answer #1

Ans:

Use binomial distribution with n=12 and p=0.2

standard deviation=sqrt(np(1-p))

=sqrt(12*0.2*(1-0.2))=1.3856